Waves are propagating on the surface of water towards the shore with a velocity of 1.5 m/s. The distance between two neighboring crests is six meters. There is a piece of wood somewhere further in the water that turns up and disappears periodically as the water waves when you are looking at it from the shore. Calculate the time interval between two turn-ups.



1
Expert's answer
2021-06-21T11:38:25-0400

"t=2\frac sv,"

"t=2\cdot \frac{6}{1.5}=8~s."